# Comprehensive Evaluation of OCT-based Automated Segmentation of Retinal Layer, Fluid and Hyper-Reflective Foci: Impact on Diabetic Retinopathy Severity Assessment

### v3

@@ II-F Semgmentation Performance Evaluation
Nikolov et al. proposed the normalized surface Dice (NSD) to estimate the deviation of surface contours within a certain threshold $\tau$(Nikolov et al., 2018). Defining a set of Euclidean distances from predicted segmentation $\hat{Y}$ to ground truth segmentation $Y$ as $\mathcal{D}_{\hat{Y}Y}$, and vice versa as $\mathcal{D}_{Y\hat{Y}}$, we obtain the subset of distances that are smaller or equal to the threshold $\tau$ as:
@@
Special attention is needed for fluid evaluation. For True Negative(TN) cases where both ground truth and predicted fluid are absent, the NSD score should be the correct prediction. For False Positive(FP) and False Negative(FN) cases where the fluid is only present in one of the ground truths or predicted segmentations, the NSD score should be zero as the incorrect prediction. We set $\tau$ to $10$ pixels for all classes, roughly $3\%$ of the shortest image edge. The model performance will be evaluated without any of the post-processing steps mentioned in the original papers.
@@ II-G Retinal Layer Thickness Analysis
Layer thickness computation was performed using the K-Nearest Neighbors (K-NN) algorithm. The layer boundaries were converted into 3D point clouds. For each data point on the upper layer, the closest corresponding point on the lower layer was identified based on Euclidean distance. The distance is properly adjusted by the voxel dimension along each axis. The thickness maps are resized to the resolution of $350\times 350$ for consistent representation. The vitreous and choroid layers were excluded from these calculations due to their unbounded nature on one side. Given the anatomical complexity of the foveal pit, the central region thickness was excluded from the analysis to ensure more reliable and interpretable measurements.
@@
Figure 2: ETDRS diagram for both left and right eyes. The diameters of the central circle, inner ring, and outer ring are 1 mm, 3 mm, and 6 mm, respectively. Nine subfields are numbered and named as follows: 1-CS(Central Subfield), 2-SI(Superior Inner), 3-NI(Nasal Inner), 4-II(Inferior Inner), 5-TI(Temporal Inner), 6-SO(Superior Outer), 7-NO(Nasal Outer), 8-IO(Inferior Outer), 9-TO(Temporal Outer).
@@ II-H Statistical Analysis
The DSC and NSD scores were calculated for model segmentation performance, and the thickness measurements were derived from the predicted segmentation. The mean DSC and NSD scores were compared within each retinal region. We used a generalized linear model (GLM) to assess the segmentation performance, thickness difference between the NPDR and PDR groups, and correlation between the thickness and visual acuity within each DR group.
Specifically, we compared the DSC and NSD scores of each pair of models with Gaussian distributions, and the performance is ranked via the effect size and p-values after the false discovery rate (FDR) correction. We compared the thickness differences between NPDR and PDR groups in each layer sector while adjusting for relevant covariates including age, gender, and duration of diabetes. Visual acuity was not included as it is considered a downstream clinical outcome rather than a demographic or biological confounder. The compound Poisson-Gamma distribution was used to model zero-inflated and highly skewed thickness measurements across DR groups while controlling for age, gender, and duration of diabetes. The correlation between the visual acuity and layer sector thickness within each DR group was modelled using Gaussian distribution while controlling for age, gender, and duration of diabetes. We converted each categorical variable to numerical values. We assigned 0 to NPDR and 1 to the PDR group, and assigned 0 to female and 1 to male. The models’ estimated coefficients (beta values) along with their 95% confidence intervals (CI) were calculated and visualized. Statistically significant results before and after FDR correction were explicitly highlighted.
@@ III-A Segmentation
Table II presents the segmentation results for four models, with values averaged across five-fold cross-validation. Tables II(a) and II(b) separately report the DSC and NSD metrics to quantify segmentation volume overlap and boundary distance respectively. SwinUNETR achieves the highest overall DSC and NSD among the evaluated models, demonstrating superior segmentation performance, particularly in the OPL, Choroid, and HRF regions. VM-UNet exhibits competitive performance, achieving the best DSC and NSD scores in the Vitreous, RNFL, and fluid regions. U-Net and SegFormer perform comparably, though U-Net slightly outperforms SegFormer in DSC across most layers, whereas SegFormer demonstrates marginally better NSD performance. These findings suggest that model predictions are less consistent in these layers, potentially due to structural complexity or segmentation challenges inherent to these regions.
@@
Figure 5(d) illustrates a representative NPDR patient with SRF. SwinUNETR and VM-UNet show superior SRF segmentation and surrounding layer boundary refinements than U-Net and SegFormer. SwinUNETR tends to over-segment both layers and fluid than VM-UNet in both layers and fluid, which is explicitly shown in the OPL layer and SRF of the fourth row.
@@
Figure 6 shows the USS and OSS for the top-2 performance models SwinUNETR and VM-UNet. VM-UNet shows lower USS than SwinUNETR in most regions except for HRF, and it has less over-segmentation in most areas except for INL, OPL, EZ and OS. Overall, using the 0.2 cutoff value, both models tend to under-segment in OPL, EZ and OS layers, plus the fluid and HRF. Over-segmentation is observed in the INL, OPL, ONL+IS, EZ and OS layers.
@@ III-B Thickness
Fig 7 shows the violin plot about the distribution of the ground-truth-derived-thickness across different retinal layers and sectors for NPDR and PDR groups. The outliers are removed outside the $5^{th}$ and $95^{th}$ percentiles, allowing a more robust interpretation of group differences. The diamond markers in each subplot show the mean thickness of each DR group without outlier removal. The PDR group has larger mean values and broader distributions in most sectors of fluid and HRF, whereas the NPDR group has larger mean thickness in most sectors of GCL+IPL and all sectors of ONL+IS.
Figure 7: The distribution of layer thickness and fluid volume measurements across different sectors for patients diagnosed with NPDR and PDR. The outliers were removed outside the $5^{th}$ and $95^{th}$ percentiles. The diamond markers represent the mean thickness for each group, calculated from the original data without outlier removal.
@@
Figure 8: Coefficient plot showing the statistical difference of thickness measurements between the model predicted segmentation and ground truth based on GLM. The horizontal dashed line represents no difference in retinal thickness. The data with p-value $<$ 0.05 after FDR correction is annotated as filled markers. The data with p-value $>$ 0.05 after FDR correction but not across the reference line is marked as ’*’. Error bars represent 95% confidence intervals.
@@
Ground-truth data reveals a significantly increased thickness in the SO sector of the OPL in PDR. Additionally, significant fluid accumulation and HRF presence were observed in the SO and NI sectors, respectively. Conversely, the ONL+IS layer demonstrated significantly greater thickness in the SI and NI sectors in the NPDR group. However, after applying FDR correction, none of these differences remained statistically significant. Both segmentation models exhibited coefficient distributions consistent with the ground truth, although no statistically significant layer-sector differences were observed. Quantitative results are shown in Supplementary Table V
Figure 9: Coefficient plot illustrating the association between NPDR and PDR groups and retinal layer thickness across different retinal sectors, controlling for age, gender, and duration of diabetes. Results were derived from GLM analysis for three segmentation results: GroundTruth (blue circles), SwinUNETR (orange circles), and VM-UNet (green circles). The horizontal dashed line represents zero effect of DR diagnosis on retinal thickness. Open markers indicate non-significant associations. Markers labelled with an asterisk (*) represent data with p-values $>0.05$ after FDR correction but not across the reference line. No filled markers (statistically significant associations after FDR correction) are present in the graph. Error bars represent 95% confidence intervals.
@@
Figure 10: Coefficient plots illustrating the association between the retinal layer thickness across different retinal sectors and visual acuity (logMAR), controlling for age, gender, and duration of diabetes. The analysis was performed for NPDR shown in 10(a) and PDR shown in 10(b), respectively. Results were derived from generalized linear model regression analyses for three segmentation results: GroundTruth (blue circles), SwinUNETR (orange circles), and VM-UNet (green circles). The horizontal dashed line represents no effect of retinal thickness on visual acuity. Error bars represent 95% confidence intervals. Open markers indicate non-significant associations. Markers labelled with an asterisk (*) represent data with p-values $>0.05$ after FDR correction but not across the reference line. No filled markers (statistically significant associations after FDR correction) are present in the graph.
Figure 12 presents four examples of thickness comparisons between NPDR and PDR groups using the ETDRS diagram described in Figure 2. Each example shows four regions that are reported with significant thickness differences in Figure 9. Figures 11(a)12(b) correspond to cases of a 57-year-old male OD, a 54-year-old male OS, and a 59-year-old female OD and OS images, respectively. Each pair of patients is matched by age, gender, and eye laterality. For each retinal region, the first row displays the En Face image overlaid with the corresponding layer thickness heatmap, while the second row presents the sector-wise quantitative average thickness. The En Face image is generated using each layer’s maximum intensity projection (MIP). For fluid and HRF, the En Face projection is derived from the entire retinal body (from the ILM to the BM), with thickness representing the accumulated volume in $\mu m^{3}$. From these figures 11(a)12(b), PDR exhibits a larger thickness than NPDR in nearly all OPL sectors. Conversely, the inner sectors of ONL+IS are significantly smaller for PDR. PDR has a larger and broader distribution of fluid and HRF accumulation in most sectors than NPDR. The findings are consistent with previous results.
(a) Example of 57-year-old male patients with right eye image.
(b) Example of 54-year-old male patients with left eye image.
(a) Example of 59-year-old female patients with right eye image.
(b) Example of 59-year-old female patients with left eye image.
@@
In the NPDR group shown in Figure 10(a), ground-truth segmentation revealed that the thickening of several layers was significantly associated with worse visual acuity (higher logMAR values). These included the RNFL (SI), OPL (SO), ONL+IS (SO), and RPE (IO). VM-UNet successfully identified the significant association between OPL thickening in the SO sector and vision loss. Additionally, it reported significant associations in the SO sector of the RNFL and INL, as well as fluid accumulation in the SO sector, all correlating with reduced vision. In contrast, SwinUNETR did not detect any statistically significant associations between retinal layer thickness or pathological volumes (fluid, HRF) and visual impairment. However, after applying FDR correction, none of the associations remained statistically significant.
In the PDR group shown in Figure 10(b), the thickening of several layer sectors was significantly correlated with poorer vision, including the EZ layer (NO) and OS layer (SO, OP and TO). The thinning of the INL (II) was also significantly correlated to worse vision. Both SwinUNETR and VM-UNet correctly identified the significant association between OS thickening and vision loss in the SO and IO sectors and also highlighted a similar association in the NO sector. Furthermore, both models predicted RPE thinning in the II sector, while VM-UNet additionally captured RNFL thinning in the TO sector. However, none of these associations remained statistically significant following FDR correction.
@@ IV-A Segmentation Models Comparison
The performance differences in DSC and NSD indicate that while both models performed well, their strengths lay in different layers. VM-UNet was superior in several layers plus fluid, whereas SwinUNETR demonstrated better performance in a few layers plus HRF. U-Net and SegFormer, though competitive in some layers, exhibited weaker performance in fine layer segmentation, particularly in RNFL and OPL, where structural continuity is essential for accurate disease characterization. Although the SwinUNETR slightly outperforms VM-UNet in several regions, VM-UNet has significantly lower computational complexity (O(N)) than SwinUNETR (O($N^{2}$)), which is crucial for remote deployment in clinics with limited computational resources.
@@
Significant differences in retinal layer thickness between NPDR and PDR offer valuable insights into the progression of DR. Ground-truth analysis revealed localized OPL thickening in the SO sector in PDR, likely indicative of extracellular fluid accumulation secondary to microvascular leakage. Similarly, increased fluid volume in the SO sector and HRF burden in the NI sector align with known patterns of retinal inflammation and exudation in advanced DR patients. In contrast, the ONL+IS thickening observed in the SI and NI sectors in NPDR may represent early photoreceptor stress or compensatory swelling. Notably, both SwinUNETR and VM-UNet were able to replicate the general pattern of effect sizes seen in the ground-truth data, suggesting their suitability for detecting biologically meaningful trends despite minor segmentation discrepancies.
@@ IV-B Clinical Insights
Our findings reveal significant associations between retinal layer thickness and visual acuity within NPDR and PDR patients. In the NPDR group, ground-truth analysis revealed that thickening of the RNFL (SI), OPL (SO), ONL+IS (SO), and RPE (IO) layers was significantly associated with worse vision. These findings likely reflect early microvascular and inflammatory changes, such as localized edema progressing into irreversible neurodegeneration. VM-UNet demonstrated strong concordance with the ground-truth trends, accurately capturing the association between OPL thickening in the SO sector and reduced vision, and further identifying plausible correlations in the RNFL, INL, and fluid in clinically relevant locations. In contrast, SwinUNETR did not identify any statistically significant relationships. In the PDR group, worse vision was significantly associated with thickening in the EZ (NO) and OS layers (SO, IO, TO), as well as thinning of the INL (II). These changes may indicate cumulative photoreceptor damage and fluid accumulation in the advanced DR stage. Both segmentation models identified the association between OS thickening and vision loss in the SO and IO sectors, with additional findings in the EZ and RPE layers. The results demonstrate the potential of the models to predict major vision changes via retinal structural variations.
Despite identifying biologically plausible trends, the models demonstrated limited predictive power in establishing statistically robust associations between retinal structural changes and visual acuity. While some associations reached nominal significance, none remained significant after the FDR correction. It demonstrates the inherent challenges with limited sample sizes, high-dimensional retinal imaging data, and complex model fitting procedures for generalized linear models with multiple covariates. The lack of statistical significance does not necessarily imply an absence of true effects but rather serves as a foundation for targeted hypothesis-driven studies in larger cohorts with increased statistical power.

### v4

@@ II-F Segmentation Performance Evaluation
Nikolov et al. proposed the normalized surface Dice (NSD) to estimate the deviation of surface contours within a certain threshold $\tau$[34]. Defining a set of Euclidean distances from predicted segmentation $\hat{Y}$ to ground truth segmentation $Y$ as $\mathcal{D}_{\hat{Y}Y}$, and vice versa as $\mathcal{D}_{Y\hat{Y}}$, we obtain the subset of distances that are smaller or equal to the threshold $\tau$ as:
@@
Special attention is needed for fluid evaluation. For True Negative(TN) cases where both ground truth and predicted fluid are absent, the NSD score should be the correct prediction. For False Positive(FP) and False Negative(FN) cases where the fluid is only present in one of the ground truths or predicted segmentations, the NSD score should be zero as an incorrect prediction. We set $\tau$ to $10$ pixels for all classes, roughly $3\%$ of the shortest image edge. The model performance will be evaluated without any of the post-processing steps mentioned in the original papers.
@@ II-G Retinal Layer Thickness Analysis
Layer thickness computation was performed using the K-Nearest Neighbors (K-NN) algorithm. The layer boundaries were converted into 3D point clouds. For each data point on the upper layer, the closest corresponding point on the lower layer was identified based on Euclidean distance. The distance is properly adjusted by the voxel dimension along each axis. The thickness maps are resized to the resolution of $350\times 350$ for consistent representation. The vitreous and choroid layers were excluded from these calculations due to their unbounded nature on one side. Given that the inner retinal layers—including RNFL, GCL+IPL, INL, OPL, and ONL+IS—converge at the foveal pit, the central region was excluded from their thickness analysis to ensure more reliable and anatomically consistent measurements.
@@
Figure 2: ETDRS diagram for both left and right eyes. The diameters of the central circle, inner ring, and outer ring are 1 mm, 3 mm, and 6 mm, respectively. Nine subfields are numbered and named as follows: 1-C(Central field), 2-SI(Superior Inner), 3-NI(Nasal Inner), 4-II(Inferior Inner), 5-TI(Temporal Inner), 6-SO(Superior Outer), 7-NO(Nasal Outer), 8-IO(Inferior Outer), 9-TO(Temporal Outer).
@@ II-H Statistical Analysis
The DSC and NSD scores were calculated for model segmentation performance, and the thickness measurements were derived from the predicted segmentation. The mean DSC and NSD scores were compared within each retinal region. We used a generalized linear model (GLM) to assess the segmentation performance, the thickness difference between the NPDR and PDR groups, and the correlation between the thickness and visual acuity within each DR group.
Specifically, we compared the DSC and NSD scores of each pair of models with Gaussian distributions, and the performance is ranked via the effect size and p-values after the false discovery rate (FDR) correction. We compared the thickness differences between NPDR and PDR groups in each layer sector while adjusting for relevant covariates, including age, gender, and duration of diabetes. Visual acuity was not included as it is considered a downstream clinical outcome rather than a demographic or biological confounder. The compound Poisson-Gamma distribution was used to model zero-inflated and highly skewed thickness measurements across DR groups while controlling for age, gender, and duration of diabetes. The correlation between the visual acuity and layer sector thickness within each DR group was modelled using the Gaussian distribution while controlling for age, gender, and duration of diabetes. We converted each categorical variable to numerical values. We assigned 0 to NPDR and 1 to the PDR group, and assigned 0 to females and 1 to males. The models’ estimated coefficients (beta values) along with their 95% confidence intervals (CI) were calculated and visualized. Statistically significant results before and after FDR correction were explicitly highlighted.
@@ III-A Segmentation
Table II presents the segmentation results for four models, with values averaged across five-fold cross-validation. Tables II(a) and II(b) separately report the DSC and NSD metrics to quantify segmentation volume overlap and boundary distance, respectively. SwinUNETR achieves the highest overall DSC and NSD among the evaluated models, demonstrating superior segmentation performance, particularly in the OPL, Choroid, and HRF regions. VM-UNet exhibits competitive performance, achieving the best DSC and NSD scores in the Vitreous, RNFL, and fluid regions. U-Net and SegFormer perform comparably, though U-Net slightly outperforms SegFormer in DSC across most layers, whereas SegFormer demonstrates marginally better NSD performance. These findings suggest that model predictions are less consistent in these layers, potentially due to structural complexity or segmentation challenges inherent to these regions.
@@
Figure 5(d) illustrates a representative NPDR patient with SRF. SwinUNETR and VM-UNet demonstrate superior SRF segmentation and refined surrounding layer boundaries compared to U-Net and SegFormer. SwinUNETR tends to over-segment both layers and fluid than VM-UNet in both layers and fluid, which is explicitly shown in the OPL layer and SRF of the fourth row.
@@
Figure 6 shows the USS and OSS for the top-2 performance models SwinUNETR and VM-UNet. VM-UNet shows lower USS than SwinUNETR in most regions, except for HRF, and it has less over-segmentation in most areas, except for INL, OPL, EZ, and OS. Overall, using the 0.2 cutoff value, both models tend to under-segment in OPL, EZ and OS layers, plus the fluid and HRF. Over-segmentation is observed in the INL, OPL, ONL+IS, EZ and OS layers.
@@ III-B Thickness
Figure 7 shows the violin plot of the distribution of the ground-truth-derived thickness across different retinal layers and sectors for NPDR and PDR groups. The outliers are removed outside the $5^{th}$ and $95^{th}$ percentiles, allowing a more robust interpretation of group differences. The diamond markers in each subplot show the mean thickness of each DR group without outlier removal. The PDR group has larger mean values and broader distributions in most sectors of RNFL, fluid and HRF, whereas the NPDR group has larger mean thickness in most sectors of GCL+IPL, ONL+IS and OS.
Figure 7: The distribution of layer thickness and fluid volume measurements across different sectors for patients diagnosed with NPDR and PDR. The outliers were removed outside the $5^{th}$ and $95^{th}$ percentiles. The diamond markers represent the mean thickness for each group, calculated from the original data without outlier removal. The central region thicknesses for inner retinal layers are not shown as mentioned in Section II-G.
@@
Figure 8: Coefficient plot showing the statistical difference of thickness measurements between the model-predicted segmentation and ground truth based on GLM. The horizontal dashed line represents no difference in retinal thickness. The data with p-value $<$ 0.05 after FDR correction is annotated as filled markers. The data with p-value $>$ 0.05 after FDR correction but not across the reference line is marked as ’*’. Error bars represent 95% confidence intervals.
@@
Ground-truth data reveal a significantly increased thickness in multiple sectors (TI, SO, NO, TO) of the RNFL in PDR. Additionally, significant fluid accumulation was observed in the NI and TO sectors, respectively. Conversely, multiple sectors in GCL+IPL (SI, II, TI) and OS (SI, II, SO, NO, IO, TO) layers demonstrated significantly greater thickness in the NPDR group. After applying FDR correction, thinning in the TI sector of GCL+IPL and multiple sectors (SO, NO, IO, TO) of OS remain significant in the PDR group. Both segmentation models exhibited coefficient distributions consistent in most regions with the ground truth, especially for the significant layer sectors after FDR correction. Quantitative results are shown in Supplementary Table V
Figure 9: Coefficient plot illustrating the association between NPDR and PDR groups and retinal layer thickness across different retinal sectors, controlling for age, gender, and duration of diabetes. Results were derived from GLM analysis for three segmentation results: GroundTruth (blue circles), SwinUNETR (orange circles), and VM-UNet (green circles). The horizontal dashed line represents zero effect of DR diagnosis on retinal thickness. Open markers indicate non-significant associations. Markers labelled with an asterisk (*) represent data with p-values $>0.05$ after FDR correction but not across the reference line. Error bars represent 95% confidence intervals.
@@
Figure 10: Coefficient plots illustrating the association between the retinal layer thickness across different retinal sectors and visual acuity (logMAR), controlling for age, gender, and duration of diabetes. The analysis was performed for NPDR shown in 10(a) and PDR shown in 10(b), respectively. Results were derived from generalized linear model regression analyses for three segmentation results: GroundTruth (blue circles), SwinUNETR (orange circles), and VM-UNet (green circles). The horizontal dashed line represents no effect of retinal thickness on visual acuity. Error bars represent 95% confidence intervals. Open markers indicate non-significant associations. Markers labelled with an asterisk (*) represent data with p-values $>0.05$ after FDR correction but not across the reference line.
Figure 12 presents four examples of thickness comparisons between NPDR and PDR groups using the ETDRS diagram described in Figure 2. Each example shows four regions that are reported with significant thickness differences in Figure 9. Figures 11(a)12(b) correspond to cases of a 54-year-old female OS, a 44-year-old male OS, a 57-year-old male OD and a 53-year-old male OD, respectively. Each pair of patients is matched by age, gender, and eye laterality. For each retinal region, the first row displays the En Face image overlaid with the corresponding layer thickness heatmap, while the second row presents the sector-wise quantitative average thickness. The En Face image is generated using each layer’s maximum intensity projection (MIP). For fluid, the En Face projection is derived from the entire retinal body (from the ILM to the BM), with thickness representing the accumulated volume in $\mu m^{3}$. From these figures 11(a)12(b), PDR exhibits a larger thickness than NPDR in most RNFL sectors. Conversely, the thickness of inner sectors of GCL+IPL and most sectors of OS is significantly smaller for PDR. PDR has a larger and broader distribution of fluid accumulation in most sectors than NPDR. The findings are consistent with previous results.
(a) Example of a 54-year-old female patient with left eye image.
(b) Example of a 44-year-old male patient with left eye image.
(a) Example of a 57-year-old male patient with right eye image.
(b) Example of a 53-year-old male patient with right eye image.
@@
In the NPDR group shown in Figure 10(a), ground-truth segmentation revealed that the thickening of several layers was significantly associated with worse visual acuity (higher logMAR values). These included the RNFL (SI, NI, II, TI, SO, NO, TO), INL (SO, NO, IO), OPL (SO, IO), ONL+IS (SO), and EZ (NI, SO, NO, IO, TO). Layer thinning is significantly related to worse vision in ONL+IS (TI) and RPE (C, NI). Visual impairment was significantly correlated to fluid (C, NI, II, SSO, NO, IO) and HRF (C, SI, NI, II, TI, NO, IO) accumulation. All correlations remained significant after FDR correction. Both models predicted similar significant associations in most layer sectors, despite some false positives in OS for SwinUNETR and in RPE for VM-UNet.
In the PDR group shown in Figure 10(b), the thickening of RNFL (TO) was significantly correlated with poorer vision. The thinning of the GCL+IPL (SI) and ONL+IS (TI) was also significantly correlated with worse vision. Fluid accumulation in multiple sectors (C, SI, NI, II, TI) was significantly correlated with vision loss, but no significant signs were shown for HRF. Both SwinUNETR and VM-UNet correctly identified most significant associations in RNFL, GCL+IPL and fluid regions. However, none of these associations remained statistically significant following FDR correction.
@@ IV-A Segmentation Models Comparison
The performance differences in DSC and NSD indicate that while both models performed well, their strengths lie in different layers. VM-UNet was superior in several layers, plus fluid, whereas SwinUNETR demonstrated better performance in a few layers, plus HRF. U-Net and SegFormer, though competitive in some layers, exhibited weaker performance in fine layer segmentation, particularly in RNFL and OPL, where structural continuity is essential for accurate disease characterization. Although the SwinUNETR slightly outperforms VM-UNet in several regions, VM-UNet has significantly lower computational complexity (O(N)) than SwinUNETR (O($N^{2}$)), which is crucial for remote deployment in clinics with limited computational resources.
@@ IV-B Clinical Insights
Significant differences in retinal layer thickness between NPDR and PDR offer valuable insights into the progression of DR. Ground-truth analysis revealed localized RNFL thickening in TI, SO, NO and TO sectors, GCL+IPL thinning in the SO sector and PS thinning in SO, NO, IO and TO sectors for PDR patients. This suggests a shift in the underlying disease mechanism—from inflammation-related thickening seen in NPDR to thinning caused by neurodegeneration and ischemia in PDR. Identifying disease stages through layer-specific biomarkers could support timely intervention. Therapies targeting edema or vascular leakage may be more beneficial during NPDR, while neuroprotective approaches might be essential for PDR curation. Similarly, increased fluid volume in the NI and TO sectors aligns with known patterns of retinal inflammation and exudation in advanced DR patients. Notably, both SwinUNETR and VM-UNet were able to replicate the general pattern of effect sizes seen in the ground-truth data, suggesting their suitability for detecting biologically meaningful trends despite minor segmentation discrepancies.
Our findings reveal significant associations between retinal layer thickness and visual acuity within NPDR and PDR patients. In the NPDR group, thickening of the RNFL in multiple sectors (SI, NI, II, TI, SO, NO, TO) may indicate early axoplasmic flow disruption and localized edema due to retinal ganglion cell dysfunction or vascular leakage. Similarly, thickening in the INL (SO, NO, IO) and OPL (SO, IO) likely reflects Müller cell swelling and intraretinal fluid accumulation—both hallmarks of early diabetic retinal changes. Thickening of the ONL+IS (SO) and EZ layers (NI, SO, NO, IO, TO) could represent subretinal fluid retention, inflammatory stress, or early disorganization of photoreceptor structures. Conversely, thinning of the ONL+IS (TI) and RPE (C, NI) was also associated with visual impairment, indicating focal photoreceptor loss or RPE atrophy. These degenerative changes impair phototransduction and outer retinal support, further degrading visual acuity. Additionally, accumulation of fluid (C, NI, II, SO, NO, IO) and HRF (C, SI, NI, II, TI, NO, IO) were significantly correlated with poor vision. These fluid-related biomarkers disrupt retinal architecture and light transmission, reinforcing the clinical relevance of monitoring retinal swelling and subcellular deposits in NPDR.
In the PDR group, thickening of the RNFL in the TO sector was associated with reduced visual acuity, suggesting localized axonal swelling or edema in advanced stages. In contrast, thinning of the GCL+IPL (SI) and ONL+IS (TI) layers was also correlated with worse vision, indicative of progressive neurodegeneration and photoreceptor disruption. These findings imply the possible vascular leakage and retinal ischemia with neuronal loss. Additionally, fluid accumulation across multiple sectors (C, SI, NI, II, TI) showed significant associations with vision loss, consistent with the finding that macular edema is a major cause of visual decline. However, HRF accumulation did not demonstrate a significant association with visual acuity in this group, potentially reflecting a shift toward more diffuse or atrophic damage in late-stage disease. Nevertheless, none of these associations remained significant after FDR correction, highlighting the complex and heterogeneous retinal remodelling present in PDR.
Both segmentation models effectively captured key associations between retinal layer thickness and visual acuity in NPDR and PDR patients, demonstrating strong alignment with ground-truth findings. They reliably identified structure–function relationships in clinically relevant layers such as RNFL, GCL+IPL, ONL+IS, and fluid. However, model-derived estimates occasionally showed greater variability, with wider confidence intervals likely due to segmentation noise or structural heterogeneity in advanced disease. Despite this, the results highlight the potential of deep learning models to predict vision impairment from retinal structural changes, supporting their use in automated risk assessment and clinical decision-making.